A particle moves in the xy-plane with the position given by the parametric equations:


x = t3 - t2 - 9t +9

y = 2 - t


Find the location of the particle when = 1.


(_, _)


1
Expert's answer
2021-03-30T07:38:31-0400

"x(1) = 1^3 - 1^2 - 9 \cdot 1 + 9 = 0 \\
y(1) = 2 -1 = 1"


Answer: (0, 1)